Which of the following is an equation of a line perpendicular to y= x-1
A. 3=y-x
B. 2y-3=x
C. y=x-3
D. y+3=x
E. y=-x+3

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Which of the following is an equation of a line perpendicular to y= x-1
y= x-1 slope m = 1
Perpendicular lines have slopes that are negative reciprocals of one another...
Any line with slope m = -1 would be Perpendicular.
Selection E has slope m = -1